Square Hammer Corp. shows the following information on its 2018 income statement Sales-$244,000; Costs = $160,000; Other expenses-$7,900; Depreciation expense- $14,900; Interest expense = $14,500. Taxes-$16,345; Dividends-$11,500. In addition, youre told that the firm issued $6,000 in new equity during 2018 and redeemed $4,500 in outstanding long-term debt. a. What is the 2018 operating cash flow? (Do not round intermediate calculations.) b. What is the 2018 cash flow to creditors? (Do not round intermediate calculations.) c. What is the 2018 cash flow to stockholders? (Do not round intermediate calculations) d. If net fixed assets increased by $20,000 during the year, what was the addition to NWC? (Do not round intermediate calculations.) a Operating cash flow b. Cash flow to creditors c. Cash flow to stockholders d. Addition to NWC
Answer #1

A:

Operating cash flow = Sales- costs- other expenses- taxes

= 244000-160000-7900-16345

=59755

B: Cash flow to creditors = Interest paid+ debt redeemed

= 14500+ 4500

= 19000

C: Cash flow to stockholders= Dividends- New stock

= 11500-6000

=5500

D: Cash flow from assets = CF to creditors+ CF to stockholders

= 19000+5500

=24500

CF from assets = Operating CF – Capital expenditure- Additions to NWC

24500 = 59755- 20000- Additions to NWC

Additions to NWC= 15255